Session: Feb 1999 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

HartRAO (Jonathan Quick at Wed Feb 17 19:21:12 MET 1999) Minor failures
Local control software failure killed LO from 16:09-16:42 causing loss of data. No other known problems

Medicina (mauro at Wed Feb 17 19:37:13 MET 1999) Success
Experiment result O.K., the clear sky, formatter - gps = -79.8 usec.

Shanghai (Huang Xinyong at Wed Feb 17 20:52:50 MET 1999) Success

Shanghai (Huang Xinyong at Wed Feb 17 20:59:58 MET 1999) Success
The EXP. is OK. The Weather is cloudy.

Torun (K.Borkowski at Thu Feb 18 09:27:55 MET 1999) Success
At 13:55 UTC there was a strong wind with snowing for about 20 minutes.

Onsala (A. Polatidis at Thu Feb 18 10:51:52 MET 1999) Success

Jodrell (Alastair Gunn at Fri Feb 19 14:16:37 MET 1999) Success
No known problems.

Cambridge (Alastair Gunn at Fri Feb 19 14:17:04 MET 1999) Success
No known problems.

Effelsberg (dgraham at Mon Feb 22 15:28:42 MET 1999) Minor failures
stopped due to snow 1615-1630,1649-1730

Westerbork (Tony Foley at Mon Feb 22 21:13:39 MET 1999) Success
No known problems

Noto (C.Trigilio at Mon Feb 22 23:23:08 MET 1999) Success

Urumqi (Liu Xiang at Tue Feb 23 04:57:41 MET 1999) Success
The LO of 5GHz was repared in Shanghai, it seems work good in the session.

Noto (C.Trigilio at Fri Feb 26 11:32:45 MET 1999) Minor failures
BBC12 not locked to the right frequency.

Calibration (Jean\-Francois Desmurs at Wed Mar 3 11:14:43 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: c99c1.antab. Read the comments in the file to get further information.